# Multilingual hierarchical classification of job advertisements for job vacancy statistics

### v1

@@ 1 Introduction
A more recent machine learning approach dedicated to job offers was proposed by Turrell et al., (2022). The researchers employed a corpus of 15.2 million advertisements posted between 2008 and 2016 on Reed.co.uk for the purpose of classifying a job offer according to the 3-digit codes of the UK Office for National Statistics’ (ONS) Standard Occupational Classification (90 minor occupational groups). In order to achieve this, the researchers employed a fuzzy matching technique. The researchers report that manual assignment of occupations yielded 76% correctness (on a sample of 330 job offers), while their algorithm achieved 91% accuracy (on a sample of 67,900 job offers). However, the authors also indicate that their assessment was based on only 34% of the initially collected job offers due to the inability of the ONS algorithm to produce a confident label. In their study, European Commission, (2022) employed a pre-trained multilingual XLM-RoBERTa (Conneau et al.,, 2020) transformer model (a type of deep neural network; see, for example, Vaswani, (2017)) to map national occupational classifications to the European Skills, Competences, Qualifications and Occupations (ESCO) framework. In addition to national qualifications and the ESCO dictionary, the authors make use of descriptions drawn from the Qualification Dataset Register and European Employment Services (EURES) online job advertisements. The authors conclude that the model demonstrates promising results when applied to work in a variety of languages. Chen and Mason, (2024) put forth the utilisation of explainable AI (XAI) methodologies to develop a lexicon-based approach. The method was applied to identify traditional and green energy jobs in online advertisements, with a reported precision rate of 82%. In 2024, Eurostat, in collaboration with Cedefop, launched a text classification competition at the ISCO level. They provided a hand-coded dataset comprising 26,000 OJA, with the highest score for the lowest common ancestor at 58% (for further details, see https://statistics-awards.eu/). It should be noted that neither the data nor the algorithms are disclosed, so they cannot be used by researchers or the official statistics community.
@@
The remaining part of the paper is as follows. Section 2 describes Polish Classification of Occupations and Specialities (hereinafter KZiS) and its relation to the International Standard Classification of Occupations (ISCO). Section 3 provides theory for the proposed approach that takes into account the hierarchical structure of official classifications. Section 4 describes the dataset used for training and testing, in particular administrative data sources, hand-coded online job offers and other available sources. Section 5 contains information about the quality of the data, in particular how accurate are the experts’ coding. Section 6 presents experimental setup and its implementation. Section 7 presents results for the proposed classifiers for Polish, bilingual (Polish and English) and multilingual datasets. The paper concludes with a discussion and an exploration of the potential for extending the approach to other classifications, including the Occupational Information Network (O*NET), the European Skills, Competences, Qualifications and Occupations (ESCO) framework, and the Standard Occupational Classification (SOC). The supplementary materials provide a brief tutorial on the use of the software, as well as detailed results.
@@ 2 Classification of occupations and specializations
The KZiS is a 6-digit, hierarchically structured classification system. The hierarchy encompasses a range of occupational groups, from one-digit major groups to six-digit occupations (cf. Table 1). It develops the ISCO classification, which ends with 4-digit codes (unit groups). In this case, it is similar to the later established European Skills, Competences, Qualifications and Occupations (ESCO) classification. The KZiS classification encompasses over 2,500 occupations, which are grouped into ten major categories. In addition to the Armed Forces (group 0), the classification is primarily delineated by the level of job position (or skill) through a process of alignment with the UNESCO International Standard Classification of Education (ISCED 2013) and the levels of the Polish Qualifications Framework, as applicable to the majority of occupations within a given group.
@@ 3.1 Formal setting
For $z_{i}\in\{0,1\}$, the code and the hierarchy tree are binary. In general, the code alphabet can contain any number of symbols. Furthermore, $z_{i}$ can take values from different sets of symbols depending on the previous values in the code. In other words, the code can result in nodes of a different parity even in the same tree. While this is not a case considered in this work, the tree does not need to be perfectly balanced, resulting in code words of different lengths for different leaf nodes. We present an example of a hierarchy tree in Figure 1.
@@
Let $\mathcal{X}$ denote an instance space, and an instance $\boldsymbol{x}\in\mathcal{X}$ be associated with a single code word $\boldsymbol{z}$ corresponding to a leaf node in the hierarchy tree, which corresponds to a subset of classes $\mathcal{C}_{\boldsymbol{x}}$. being assigned to the $\boldsymbol{x}$. This subset is often called a set of relevant labels, while the compliment \$\mathcal{L}\backslash\mathcal{L}_{\boldsymbol{x}}$ is considered as irrelevant for $\boldsymbol{x}$. We identify a set $\mathcal{L}_{\boldsymbol{x}}$ of relevant labels with a binary (sparse) vector $\boldsymbol{y}=(y_{1},\ldots,y_{m})$, in which $y_{c}=1\Leftrightarrow c\in\mathcal{L}_{\boldsymbol{x}}$. By $\mathcal{Y}=\{0,1\}^{m}$ we denote a set of all possible label vectors. We assume that observations $(\boldsymbol{x},\boldsymbol{y})$ are generated independently and identically according to the probability distribution $\mathbb{P}(\boldsymbol{X}=\boldsymbol{x},\boldsymbol{Y}=\boldsymbol{y})$ (denoted later by $\mathbb{P}(\boldsymbol{x},\boldsymbol{y})$) defined on $\mathcal{X}\times\mathcal{Y}$).
@@
$L_{\ell}(\boldsymbol{h}):=\mathbb{E}_{(\boldsymbol{x},\boldsymbol{y})\sim\mathbb{P}(\boldsymbol{x},\boldsymbol{y})}\left[\ell(\boldsymbol{y},\boldsymbol{h}(\boldsymbol{x}))\right]\,,$ (1)
where $\ell(\boldsymbol{y},\hat{\boldsymbol{y}})$ is the (task) loss.
In particular, as our goal is to use the classifier for population analysis, we aim at estimating the conditional probabilities of classes $\eta_{c}(\boldsymbol{x}):=\mathbb{P}(y_{c}=1|\boldsymbol{x})=\mathbb{P}(\boldsymbol{z}^{l}=\boldsymbol{z}^{l}(c)|\boldsymbol{x})$. To obtain the estimates of conditional probabilities, one can use the class-wise log loss as a surrogate objective:
@@ 3.2 Modeling probabilities under class hierarchy
We are not only interested in possible accurate estimates of conditional probabilities $\eta_{c}(\boldsymbol{x})$ for each class $c$, we additionally require for the predictions of the classifier to be coherent under class hierarchy, that is, obey the relations presented in Equations 6 and 5. Naively training the classifier using class-wise log loss 2 does not guarantee that class relations will hold. Because of that, in this work, we consider two classical approaches that guarantee coherency of prediction under class hierarchy.
We refer to the first approach as a “bottom-up” (cf. Barbedo and Lopes,, 2006). The bottom-up approach is one of the simplest to deal with the hierarchical classification problem. In this type of classifier, the conditional probability estimator of classes is trained only for classes in the leaf nodes of the hierarchy tree (i.e., 6-digit KZiS codes), and information about the class hierarchy is ignored. Then, during inference, after predicting probabilities for leaf labels, the predictions for labels higher in the hierarchy are reconstructed by summing the estimates of their direct children as in 5 until the root node is reached.
@@
The classifier then is formed by training probability estimator for every $\mathbb{P}(z_{i}(c)|\boldsymbol{z}^{i-1}(c),\boldsymbol{x})$. This can be done using the same log loss 2. By using logistic loss and a linear model $f_{\boldsymbol{z}^{i}}(\boldsymbol{x})$ in each node $\boldsymbol{z}^{i}$ for estimating $\mathbf{P}\left(z_{i}\mid\boldsymbol{z}^{i-1},\boldsymbol{x}\right)$, we obtain the popular formulation of HSM. Let us notice that since we deal here with a multi-class distribution, we have that:
$\sum_{c}\mathbb{P}\left(z_{i}=c\mid\boldsymbol{z}^{i-1},\boldsymbol{x}\right)=1$ (8)
This approach is known in the literature as the nested dichotomies (Fox,, 1997) in statistics, hierarchical softmax approach (Morin and Bengio,, 2005), conditional probability estimation trees (Beygelzimer et al.,, 2009) or probabilistic classifier trees (Dembczynski et al.,, 2016)
@@ 3.3 Statistical challenge of long-tail distribution
In our task, the classifier must predict a set of classes out of 2,911 possible ones. Figure 2 illustrates the long-tail distribution present in the dataset, which is described in detail in Section 4. It is evident that almost one-third of the classes are represented by less than 10 examples, while only around 250 classes have more than 100.
@@ 3.4 Proposed model architecture
To address long-tail distribution, instead of training the classifier from scratch, we opted for using a pre-trained transformer model. We are focusing on the popular family of models based on the Bidirectional Encoder Representations from Transformers (BERT; Devlin, (2018)) architecture.
@@ 4.1 Official classifications and definitions
  • Job description search engine77endnote: 7Available at https://psz.praca.gov.pl/rynek-pracy/bazy-danych/klasyfikacja-zawodow-i-specjalnosci/wyszukiwarka-opisow-zawodow (as of 2024.07.18). – This service provides a comprehensive description of all occupations, including the name, code, synthesis, and a list of elementary and additional job tasks. For illustrative purposes, please refer to Figure 3. It is important to note that the descriptions of the occupations exhibit considerable variation in terms of both detail and completeness.
  • INFOdoradca+ webservice88endnote: 8https://psz.praca.gov.pl/rynek-pracy/bazy-danych/infodoradca (as of 2024.07.18). – The service comprises a curated list of 1,000 occupations, offering a comprehensive overview of the profession or job role in question. It provides detailed information on the specific skills, qualifications and authorisations required for entry into the profession or job role, as well as the tasks and competencies necessary for success in the field. For an illustrative example, please refer to Figure 3, which depicts a description from the INFOdoradca+ web service.
@@
In order to ensure the accuracy of online job advertisement classification, it is essential to utilise a comprehensive range of resources beyond official dictionaries or a limited number of advertisements from the Civil Service. Consequently, we have elected to employ the actual job postings from two primary sources: the Central Job Offers Database (CBOP), which encompasses all job advertisements submitted to Public Employment Offices (PEOs), and hand-coded job advertisements from a multitude of online sources. The subsequent subsections will delineate the characteristics of these datasets.
4.2 Central Job Offers Database
Public employment services in Poland include public employment offices (PEOs), which operate at NUTS4 level (Nomenclature des unités territoriales statistiques, Nomenclature of Territorial Units for Statistics) and are responsible for registering and managing unemployment. This means that there are two sources of data about employment in Poland: (i) the Labor Force Survey (LFS) and (ii) registered data collected by PEOs. The register maintained by PEOs contains detailed characteristics of unemployed persons and job offers. The number of registered unemployed may be higher than the LFS estimate because registration is connected with free health insurance and unemployment benefits.
@@ 4.2 Central Job Offers Database
Figure 4: Example job posting from CBOP
@@
Job offers may be submitted via a paper or online form, following a highly structured format1212endnote: 12For an illustrative example, please refer to this website: https://warszawa.praca.gov.pl/zgloszenie-oferty-pracy (in Polish). In Supplementary Materials we provide an example paper form in Polish and English translated using DeepL software.. The requisite information encompasses the entity’s name, address, and REGON/NIP identifier; the job title; the number of vacancies; the type of contract; a description of the responsibilities and requirements for the applicants. An example of a job posting is provided in Figure 4. It is important to note that the option to specify the job occupation according to the official classification is available on the online form, though this is not a mandatory requirement. In accordance with the findings of our discourse with PEOs, it has been established that the majority of employers refrain from populating this field, and instead, PEOs personnel undertake the coding of job advertisements according to the official KZiS classification manually. Regrettably, we are unable to present any numerical results to substantiate our claims, as this information is not accessible through CBOP database. In Section 5, we examine the quality of CBOP coding using a sample of job ads. Information about the occupation is policy-relevant. The Ministry of National Education annually announces the forecast of the demand for employees in vocational education occupations on the national and regional labour markets1313endnote: 13See https://www.gov.pl/web/edukacja/prognoza-zapotrzebowania-na-pracownikow-w-zawodach-szkolnictwa-branzowego-na-krajowym-i-wojewodzkim-rynku-pracy-2024. Based on the forecast the Ministry differentiates the educational part of the general subsidy for vocational education students between local government units. The regional governments consider the forecast in preparing an annual list of occupations for which the costs of vocational training for young employees is reimbursed. The forecast is also used by regional labour market councils in issuing opinions on the validity of education in occupations.
@@
The CBOP data is accessible via the API for registered users1414endnote: 14More information is available at the following website https://oferty.praca.gov.pl/portal/index.cbop#/dlaInt. Each job advertisement comprises 174 fields in JSON format, including an occupation code according to the KZiS and descriptions provided by the employer or recruitment agency. The data used in this study was downloaded on a daily basis from the beginning of 2022 to the end of 2023. The data was then subjected to a process of cleaning, whereby erroneous codes or descriptions with missing data in all fields were removed1515endnote: 15The codes used for this purpose are available for download online at https://github.com/OJALAB/CBOP-datasets. Following this, the data set comprised over 822,000 ads with 2,468 unique occupation codes and job ads descriptions, based on the aforementioned criteria.
@@
The distribution of KZiS codes was markedly uneven, with 145 codes having only a single example, 128 codes having two examples, and codes with over 5,000 examples1616endnote: 16The five most prevalent codes were: The following job titles were identified: 522301 ”Sprzedawca” (Sales Assistant), 432103 ”Magazynier” (Warehouse Operator), 931301 ”Pomocniczy robotnik budowlany” (Construction Worker’s Helper), 911207 ”Pracownik utrzymania czystości” (Cleaner), 515303 ”Robotnik gospodarczy” (Maintenance Worker).. To reduce the size of CBOP database, a stratified sample of the CBOP dataset was created, with strata defined by KZiS code and the number of characters in the description. A total of 11,572 strata were formed. Table 2 presents information regarding the number of advertisements by the number of characters in the advertisement description.
Table 2: Number of CBOP ads by the number of characters in the description
@@
$>500$ 198,614
[…]

### v2

@@ 1 Introduction
A more recent machine learning approach dedicated to job offers was proposed by Turrell et al., (2022). The researchers employed a corpus of 15.2 million advertisements posted between 2008 and 2016 on Reed.co.uk for the purpose of classifying a job offer according to the 3-digit codes of the UK Office for National Statistics’ (ONS) Standard Occupational Classification (90 minor occupational groups). In order to achieve this, the researchers employed a fuzzy matching technique. The researchers report that manual assignment of occupations yielded 76% correctness (on a sample of 330 job offers), while their algorithm achieved 91% accuracy (on a sample of 67,900 job offers). However, the authors also indicate that their assessment was based on only 34% of the initially collected job offers due to the inability of the ONS algorithm to produce a confident label. In their study, European Commission, (2022) employed a pre-trained multilingual XLM-RoBERTa (Conneau et al.,, 2020) transformer model (a type of deep neural network; see, for example, Vaswani et al., (2017)) to map national occupational classifications to the European Skills, Competences, Qualifications and Occupations (ESCO) framework. In addition to national qualifications and the ESCO dictionary, the authors make use of descriptions drawn from the Qualification Dataset Register and European Employment Services (EURES) online job advertisements. The authors conclude that the model demonstrates promising results when applied to work in a variety of languages. Chen and Mason, (2024) put forth the utilisation of explainable AI (XAI) methodologies to develop a lexicon-based approach. The method was applied to identify traditional and green energy jobs in online advertisements, with a reported precision rate of 82%. In 2024, Eurostat, in collaboration with Cedefop, launched a text classification competition at the ISCO level. They provided a hand-coded dataset comprising 26,000 OJA, with the highest score for the lowest common ancestor at 58% (for further details, see https://statistics-awards.eu/). It should be noted that neither the data nor the algorithms are disclosed, so they cannot be used by researchers or the official statistics community.
@@
The remaining part of the paper is as follows. Section˜2 describes Polish Classification of Occupations and Specialities (hereinafter KZiS) and its relation to the International Standard Classification of Occupations (ISCO). Section˜3 provides theory for the proposed approach that takes into account the hierarchical structure of official classifications. Section˜4 describes the dataset used for training and testing, in particular administrative data sources, hand-coded online job offers and other available sources. Section˜5 contains information about the quality of the data, in particular how accurate are the experts’ coding. Section˜6 presents experimental setup and its implementation. Section˜7 presents results for the proposed classifiers for Polish, bilingual (Polish and English) and multilingual datasets. The paper concludes with a discussion and an exploration of the potential for extending the approach to other classifications, including the Occupational Information Network (O*NET), the European Skills, Competences, Qualifications and Occupations (ESCO) framework, and the Standard Occupational Classification (SOC). The supplementary materials provide a brief tutorial on the use of the software, as well as detailed results.
@@ 2 Classification of occupations and specializations
The KZiS is a 6-digit, hierarchically structured classification system. The hierarchy encompasses a range of occupational groups, from one-digit major groups to six-digit occupations (cf. Table˜1). It develops the ISCO classification, which ends with 4-digit codes (unit groups). In this case, it is similar to the later established European Skills, Competences, Qualifications and Occupations (ESCO) classification. The KZiS classification encompasses over 2,500 occupations, which are grouped into ten major categories. In addition to the Armed Forces (group 0), the classification is primarily delineated by the level of job position (or skill) through a process of alignment with the UNESCO International Standard Classification of Education (ISCED 2013) and the levels of the Polish Qualifications Framework, as applicable to the majority of occupations within a given group.
@@ 3.1 Formal setting
For $z_{i}(c)\in\{0,1\}$, the code and the hierarchy tree are binary. In general, the code alphabet can contain any number of symbols. Furthermore, $z_{i}$ can take values from different sets of symbols depending on the previous values in the code. In other words, the code can result in nodes of a different parity even in the same tree. While this is not a case considered in this work, the tree does not need to be perfectly balanced, resulting in code words of different lengths for different leaf nodes. We present an example of a hierarchy tree in Figure˜1.
@@
Let $\mathcal{X}$ denote an instance space, and an instance $\boldsymbol{x}\in\mathcal{X}$ be associated with a single code word $\boldsymbol{z}$ corresponding to a leaf node in the hierarchy tree, which corresponds to a subset of classes $\mathcal{C}_{\boldsymbol{x}}$. being assigned to the $\boldsymbol{x}$. This subset is often called a set of relevant labels, while the compliment $\mathcal{L}\backslash\mathcal{L}_{\boldsymbol{x}}$ is considered as irrelevant for $\boldsymbol{x}$. We identify a set $\mathcal{L}_{\boldsymbol{x}}$ of relevant labels with a binary (sparse) vector $\boldsymbol{y}=(y_{1},\ldots,y_{m})$, in which $y_{c}=1\Leftrightarrow c\in\mathcal{L}_{\boldsymbol{x}}$. By $\mathcal{Y}=\{0,1\}^{m}$ we denote a set of all possible label vectors. We assume that observations $(\boldsymbol{x},\boldsymbol{y})$ are generated independently and identically according to the probability distribution $\mathbb{P}(\boldsymbol{X}=\boldsymbol{x},\boldsymbol{Y}=\boldsymbol{y})$ (denoted later by $\mathbb{P}(\boldsymbol{x},\boldsymbol{y})$) defined on $\mathcal{X}\times\mathcal{Y}$).
@@
$L_{\ell}(\boldsymbol{h})\coloneqq\mathbb{E}_{(\boldsymbol{x},\boldsymbol{y})\sim\mathbb{P}(\boldsymbol{x},\boldsymbol{y})}\left[\ell(\boldsymbol{y},\boldsymbol{h}(\boldsymbol{x}))\right]\,$, (1)
[…]
In particular, as our goal is to use the classifier for population analysis, we aim at estimating the conditional probabilities of classes $\eta_{c}(\boldsymbol{x})\coloneqq\mathbb{P}(y_{c}=1|\boldsymbol{x})=\mathbb{P}(\boldsymbol{z}^{l}=\boldsymbol{z}^{l}(c)|\boldsymbol{x})$. To obtain the estimates of conditional probabilities, one can use the class-wise log loss as a surrogate objective:
@@ 3.2 Modeling probabilities under class hierarchy
We are not only interested in possible accurate estimates of conditional probabilities $\eta_{c}(\boldsymbol{x})$ for each class $c$, we additionally require for the predictions of the classifier to be coherent under class hierarchy, that is, obey the relations presented in Equations˜6 and 5. Naively training the classifier using class-wise log loss 2 does not guarantee that class relations will hold. Because of that, in this work, we consider two classical approaches that guarantee coherency of prediction under class hierarchy.
We refer to the first approach as a “bottom-up” (cf. Barbedo and Lopes,, 2007). The bottom-up approach is one of the simplest to deal with the hierarchical classification problem. In this type of classifier, the conditional probability estimator of classes is trained only for classes in the leaf nodes of the hierarchy tree (i.e., 6-digit KZiS codes), and information about the class hierarchy is ignored. Because, out of these classes, only one can be positive at for instance, one can use categorical cross-entropy loss (a generalization of log loss) to train the probability estimator:
$\ell_{\text{cce}}(\boldsymbol{y},\boldsymbol{h}(\boldsymbol{x}))=\sum_{c\in\mathcal{C}^{\prime}}y_{c}\log(h_{c}(\boldsymbol{x}))\,$, (7)
where $\mathcal{C}^{\prime}$ is some subset of all the classes $\mathcal{C}$, in this case being a subset of classes without any children. During inference, after predicting probabilities for leaf labels, the predictions for labels higher in the hierarchy are reconstructed by summing the estimates of their direct children as in 5 until the root node is reached.
@@
Let us notice that since we deal here with a multi-class distribution, we have that:
$\sum_{z_{i}}\mathbb{P}\left(z_{i}\mid\boldsymbol{z}^{i-1},\boldsymbol{x}\right)=1$ (9)
The classifier is then formed by the training probability estimator for every $\mathbb{P}(z_{i}(c)|\boldsymbol{z}^{i-1}(c),\boldsymbol{x})$. This can be done using the cross-entropy loss 7 introduced earlier. This approach is known in the literature as the nested dichotomies (Fox,, 1997) in statistics, hierarchical softmax (HSM) approach (Morin and Bengio,, 2005), conditional probability estimation trees (Beygelzimer et al.,, 2009) or probabilistic classifier trees (Dembczyński et al.,, 2016)
@@ 3.3 Statistical challenge of long-tail distribution
In our task, the classifier must predict a set of classes out of 2,911 possible ones. Figure˜2 illustrates the long-tail distribution present in the dataset, which is described in detail in Section˜4. It is evident that almost one-third of the classes are represented by less than 10 examples, while only around 250 classes have more than 100.
@@ 3.4 Proposed model architecture
To address long-tail distribution, instead of training the classifier from scratch, we opted for using a pre-trained transformer model. We are focusing on the popular family of models based on the Bidirectional Encoder Representations from Transformers (BERT; Devlin et al., (2019)) architecture.
@@ 4.1 Official classifications and definitions
  • Job description search engine77endnote: 7Available at https://psz.praca.gov.pl/rynek-pracy/bazy-danych/klasyfikacja-zawodow-i-specjalnosci/wyszukiwarka-opisow-zawodow (as of 2024.07.18). – This service provides a comprehensive description of all occupations, including the name, code, synthesis, and a list of elementary and additional job tasks. For illustrative purposes, please refer to Figure˜3. It is important to note that the descriptions of the occupations exhibit considerable variation in terms of both detail and completeness.
  • INFOdoradca+ webservice88endnote: 8https://psz.praca.gov.pl/rynek-pracy/bazy-danych/infodoradca (as of 2024.07.18). – The service comprises a curated list of 1,000 occupations, offering a comprehensive overview of the profession or job role in question. It provides detailed information on the specific skills, qualifications and authorisations required for entry into the profession or job role, as well as the tasks and competencies necessary for success in the field. For an illustrative example, please refer to Figure˜3, which depicts a description from the INFOdoradca+ web service.
@@
In order to ensure the accuracy of online job advertisement classification, it is essential to utilise a comprehensive range of resources beyond official dictionaries or a limited number of advertisements from the Civil Service. Consequently, we have chosen to employ actual job postings from two primary sources: the Central Job Offers Database (hereinafter ePraca; previously known as CBOP), which encompasses all job advertisements submitted to Public Employment Offices (PEOs), and hand-coded job advertisements from multiple online sources. The subsequent subsections will delineate the characteristics of these datasets.
4.2 Central Job Offers Database – ePraca
Public employment services in Poland include public employment offices (PEOs), which operate at LAU1 (Local Administrative Units) level and are responsible for registering and managing unemployment. This means that there are two sources of data about employment in Poland: (i) the Labor Force Survey (LFS) and (ii) registered data collected by PEOs. The register maintained by PEOs contains detailed characteristics of unemployed persons and job offers. The number of registered unemployed may be higher than the LFS estimate because registration is connected with free health insurance and unemployment benefits.
@@ 4.2 Central Job Offers Database – ePraca
Figure 4: Example job posting from ePraca
@@
Job offers may be submitted via a paper or online form, following a highly structured format1212endnote: 12For an illustrative example, please refer to this website: https://warszawa.praca.gov.pl/zgloszenie-oferty-pracy (in Polish). In Supplementary Materials we provide an example paper form in Polish and English translated using DeepL software.. The requisite information encompasses the entity’s name, address, and REGON/NIP identifier; the job title; the number of vacancies; the type of contract; a description of the responsibilities and requirements for the applicants. An example of a job posting is provided in Figure˜4. It is important to note that the option to specify the job occupation according to the official classification is available on the online form, though this is not a mandatory requirement. In accordance with the findings of our discourse with PEOs, it has been established that the majority of employers refrain from populating this field, and instead, PEOs personnel undertake the coding of job advertisements according to the official KZiS classification manually. Regrettably, we are unable to present any numerical results to substantiate our claims, as this information is not accessible through ePraca database. In Section˜5, we examine the quality of ePraca coding using a sample of job ads. Information about the occupation is policy-relevant. The Ministry of National Education annually announces the forecast of the demand for employees in vocational education occupations on the national and regional labour markets1313endnote: 13See https://www.gov.pl/web/edukacja/prognoza-zapotrzebowania-na-pracownikow-w-zawodach-szkolnictwa-branzowego-na-krajowym-i-wojewodzkim-rynku-pracy-2024. Based on the forecast the Ministry differentiates the educational part of the general subsidy for vocational education students between local government units. The regional governments consider the forecast in preparing an annual list of occupations for which the costs of vocational training for young employees is reimbursed. The forecast is also used by regional labour market councils in issuing opinions on the validity of education in occupations.
@@
The ePraca data is accessible via the API for registered users1414endnote: 14More information is available at the following website https://oferty.praca.gov.pl/portal/index.cbop#/dlaInt. Each job advertisement comprises 174 fields in JSON format, including an occupation code according to the KZiS and descriptions provided by the employer or recruitment agency. The data used in this study was downloaded on a daily basis from the beginning of 2022 to the end of 2023. The data was then subjected to a process of cleaning, whereby erroneous codes or descriptions with missing data in all fields were removed1515endnote: 15The codes used for this purpose are available for download online at https://github.com/OJALAB/CBOP-datasets. Following this, the data set comprised over 822,000 fully labelled ads with 2,468 unique occupation codes and job ads descriptions, based on the aforementioned criteria.
@@
The distribution of KZiS codes was markedly uneven, with 145 codes having only a single example, 128 codes having two examples, and codes with over 5,000 examples1616endnote: 16The five most prevalent codes were: The following job titles were identified: 522301 ”Sprzedawca” (Sales Assistant), 432103 ”Magazynier” (Warehouse Operator), 931301 ”Pomocniczy robotnik budowlany” (Construction Worker’s Helper), 911207 ”Pracownik utrzymania czystości” (Cleaner), 515303 ”Robotnik gospodarczy” (Maintenance Worker).. To reduce the size of the ePraca database and speed up training, a stratified sample of the ePraca dataset was created, with strata defined by KZiS code and the number of characters in the description. A total of 11,572 strata were formed. Table˜2 presents information regarding the number of advertisements by the number of characters in the advertisement description.
Table 2: Number of ePraca ads by the number of characters in the description
@@
>500>500 198,614
From each stratum, a simple random sample was selected, with the size determined as $\max\left\{1,\lfloor 0.2\times N_{h}\rceil\right\}$, where $N_{h}$ is the strata sample size and $\lfloor\rceil$ denotes rounding to the nearest integer. The final sample size was 167,244 (around 20% of the ePraca dataset).